solve the following equation
3/(z-3) + 5/(z-1) = 8/(z-2)

Answer by perfectdose(7)


First we multiply by our three different denominators (z-3),(z-1),(z-2) to clear fractions.
3(z-1)(z-2)+5(z-2)(z-3)=8(z-1)(z-3)
now time to foil
3(z^2-1x-2x+2)+5(z^2-2z-3z+6)=8(z^2-1z-3z+3)
combine like terms
3(z^2-3z+2)+5(z^2-5z+6)=8(z^2-4z+3)
Distribute the number to each term in the parentheses.
3z^2-9z+6+5z^2-25z+30=8z^2-32z+24
Combine like terms
8z^2-34z+36=8z^2-32z+24
Solve
12 = 2z
6 = z
:)